OpenStudy (anonymous):

√2x =8 solve for x ***** metal ****

OpenStudy (johnweldon1993):

\[\sqrt{2x} = 8\] We want to get rid of that square root....so lets square both sides \[\sqrt{2x}^2 = 8^2\] This now becomes \[2x = 64\] Now divide both sides by 2 to solve for 'x' \[x = 32\]